in a basketball game, adam made 19 baskets, some 2-point shots and some 3-points shots, for a total of 40 point. how many 3-points baskets did he make
Answer by jaydducote(11)

You have to make a system of equations.
Let's make x=number of two pointers and y=number of threes
We know he made 19 shots, so x+y=19
We also know he scored 40 points, so 2X+3y=40. We know this because of two points and three points for each type of basket. Nowe we need to find a way to combine equations. Since the problem asks for how man 3-point baskets he made, we want to solve for y. This means that we first need to get x in terms of y. Take the first equation and solve for x;
x+y=19
subtract y from both sides
x=19-y
now substitute that value for x in the second equation:
2x+3y=40
2(19-y)+3y=40
distribute the 2
38-2y+3y=40
combine like terms
38+y=40
subtract 38 from both sides
y=2
Adam made 2 three-point baskets.